Qs: 04 Let T(x, y, z, w) = ( 3y -4z, x +2y+4z-w, 7z, -y-w)

Write the standard matrix for T. Check if T is one-to-one and onto.
1
Expert's answer
2020-06-15T15:30:37-0400

"T(x, y, z, w) = ( 3y -4z, x +2y+4z-w, 7z, -y-w)"

Write the standard matrix for T.


"A=\begin{bmatrix}
 0 & 3 & -4 & 0\\
 1 & 2 & 4 & -1 \\
 0 & 0 & 7 & 0\\
 0 & -1 & 0 & -1
\end{bmatrix}"

One-to-one is the same as onto for square matrices.

In general, a transformation T is both one-to-one and onto if and only if "T(\bf x )=\bf b"has exactly one solution for all "\bf b" in "\Bbb{R}^m."


"\begin{bmatrix}
 0 & 3 & -4 & 0 & &0\\
 1 & 2 & 4 & -1 & &0 \\
 0 & 0 & 7 & 0 & &0\\
 0 & -1 & 0 & -1 & &0
\end{bmatrix}"

Swap rows 1 and 2:


"\begin{bmatrix}
 1 & 2 & 4 &-1& &0\\
 0 & 3 & -4 & 0 & &0 \\
 0 & 0 & 7 & 0 & &0\\
 0 & -1 & 0 & -1 & &0
\end{bmatrix}"

"R_2={R_2\/3}"


"\begin{bmatrix}
 1 & 2 & 4 &-1& &0\\
 0 & 1 & -4/3 & 0 & &0 \\
 0 & 0 & 7 & 0 & &0\\
 0 & -1 & 0 & -1 & &0
\end{bmatrix}"

"R_1=R_1-(2)R_2"


"\begin{bmatrix}
 1 & 0 & 20/3 &-1& &0\\
 0 & 1 & -4/3 & 0 & &0 \\
 0 & 0 & 7 & 0 & &0\\
 0 & -1 & 0 & -1 & &0
\end{bmatrix}"

"R_4=R_4+R_2"


"\begin{bmatrix}
 1 & 0 & 20/3 &-1& &0\\
 0 & 1 & -4/3 & 0 & &0 \\
 0 & 0 & 7 & 0 & &0\\
 0 & 0 & -4/3 & -1 & &0
\end{bmatrix}"

"R_3=R_3\/7"


"\begin{bmatrix}
 1 & 0 & 20/3 &-1& &0\\
 0 & 1 & -4/3 & 0 & &0 \\
 0 & 0 & 1 & 0 & &0\\
 0 & 0 & -4/3 & -1 & &0
\end{bmatrix}"

"R_1=R_1-(20\/3)R_3"


"\begin{bmatrix}
 1 & 0 & 0 &-1& &0\\
 0 & 1 & -4/3 & 0 & &0 \\
 0 & 0 & 1 & 0 & &0\\
 0 & 0 & -4/3 & -1 & &0
\end{bmatrix}"

"R_2=R_2+(4\/3)R_3"


"\begin{bmatrix}
 1 & 0 & 0 &-1& &0\\
 0 & 1 & 0 & 0 & &0 \\
 0 & 0 & 1 & 0 & &0\\
 0 & 0 & -4/3 & -1 & &0
\end{bmatrix}"

"R_4=R_4+(4\/3)R_3"


"\begin{bmatrix}
 1 & 0 & 0 &-1& &0\\
 0 & 1 & 0 & 0 & &0 \\
 0 & 0 & 1 & 0 & &0\\
 0 & 0 & 0 & -1 & &0
\end{bmatrix}"

"R_1=R_1-R_4"


"\begin{bmatrix}
 1 & 0 & 0 & 0& &0\\
 0 & 1 & 0 & 0 & &0 \\
 0 & 0 & 1 & 0 & &0\\
 0 & 0 & 0 & -1 & &0
\end{bmatrix}"

"R_4=R_4\cdot(-1)"


"\begin{bmatrix}
 1 & 0 & 0 & 0& &0\\
 0 & 1 & 0 & 0 & &0 \\
 0 & 0 & 1 & 0 & &0\\
 0 & 0 & 0 & 1 & &0
\end{bmatrix}"

The columns of matrix are linearly independent, which happens precisely when the matrix has a pivot position in every column. 

Therefore "T" is one-to-one.